The Candy Emporium sells two types of hard candy--type A sells for $1.50 per pound and type B sells for $2.00 per pound. How much of each type of candy should be combined to obtain 20 pounds of a mixture which sells for $1.80 per pound?

Answer by mananth(16946)

price ----------------quantity
Candy A 1.5 ----------------x pounds
Candy B 2 ----------------20-x pounds
Mixture 1.80% ----------------20

1.5x+2(20-x)=1.8*20

1.5x+40-2x=36
1.5x-2x=36-40
-0.5x=-4
/-0.5
x=8 pounds $1.50
12 pounds $2.00